OpenStudy (anonymous):

f(x)= x^2 / sqrt(x^2-a^2) , find f'(x)

OpenStudy (agreene):

\[\frac{\delta}{\delta x}[\frac{x^2}{\sqrt{x^2-a^2}}]=\frac{x^3-a^2x}{(x^2-a^2)^{3/2}}\]

OpenStudy (anonymous):

Can you show steps?

OpenStudy (agreene):

Product Rule: \[=x^2(\frac{d}{dx}(\frac{1}{\sqrt{x^2-a^2}})+\frac{2x}{\sqrt{x^2-a^2}}\] Chain Rule: \[\large =x^2-\frac{\frac{d}{dx}(x^2-a^2)}{2(x^2-a^2)^{3/2}}+\frac{2x}{\sqrt{x^2-a^2}}\] \[\large=\frac{2x}{\sqrt{x^2-a^2}}-\frac{x^2(2x)}{2(x^2-a^2)^{3/2}}\] do the subtraction and you get what i first put.

OpenStudy (anonymous):

Thanks!